Grade 12th passPhysical Chemistry

how 1 gram mole eguals to 1 mole.?
12 gram of carbon contains 1 mole of atoms.
what is 1 gram mole

famy francis
9 Years agoGrade 12th pass

1 Answer

Vaibhav Singh
9 Years ago
It is no way related to mass.1amu = 1/12 of mass of C12 isotope and 1 amu = 1/6.023x10^23 g.Therefore, 1 g = 1amu x 1 mole(Here mole represents only number ie 6.023x10^23 not any physical quantity) We write 1 g atom to mean 1 Mole of atoms. Similarly we write 1 g ion or molecule to refer to 1 Mole ions or molecules respectively.